### v1

# Reconstructing a Heavy Neutral Lepton at the LHC

Abstract
[…]
Reconstructing a Heavy Neutral Lepton at the LHC
Pablo de la Torre, Manuel Masip, Fuensanta Vilches
Departamento de Física Teórica y del Cosmos
Universidad de Granada, E-18071 Granada, Spain
pdelatorre@ugr.es      masip@ugr.es      fuenvilches@ugr.es
1 Introduction
Heavy fermion singlets with no electric charge provide the simplest UV completion of the Standard Model (SM) that is able to generate the dim-5 Weinberg operator [1], which arguably makes them the most motivated scenario for BSM physics. Let us briefly discuss the minimal setup (see the generic discussion in [2], the majoron model in [3] or the dark matter model in [4]) using 2-component spinors***In this notation $\chi$ is a left-handed spinor whereas its conjugate contravariant $\bar{\chi}$ is right handed (we omit Lorentz indexes). For example, to define the electron we need two of these left handed spinors with opposite charge: $e$ and $e^{c}$. A mass term would then read $m\left(ee^{c}+\bar{e}\bar{e^{c}}\right)=m\,\bar{\Psi}_{e}\Psi_{e}$, where $\Psi_{e}$ is the 4-component electron in the chiral representation of $\gamma^{\mu}$: $\Psi_{e}=\begin{pmatrix}e\\ \bar{e^{c}}\end{pmatrix}$ and $\bar{\Psi}_{e}=\left(e^{c}\;\bar{e}\right)$..
To explain neutrino masses we need at least two heavy neutral leptons (HNLs): two bi-spinors $(N,N^{c})$ of opposite lepton number $L=\pm 1$ combined into a Dirac field of mass $M$. Assuming no extra Higgses, lepton-number conservation and only dim $\leq 4$ operators we have
$-{\cal L}\supset M\,NN^{c}+y_{\nu}\,HLN^{c}+{\rm h.c.}\,,$ (1)
where $H=(h^{+}\;h^{0})$, $L=(\nu\;\ell)$ and, with all generality, we have redefined the 3 lepton doublets to find the flavor combination in the Yukawa. After electroweak (EW) symmetry breaking the neutrino mass matrix reads
${\cal M}=\begin{pmatrix}&&&\cdot&0\\ &0&&\cdot&0\\ &&&\cdot&m\\ \cdot&\cdot&\cdot&\cdot&M\\ 0&0&m&M&\cdot\end{pmatrix},$ (2)
where $m=y_{\nu}v/\sqrt{2}$ and the dots indicate terms forbidden by $L$. This rank-2 matrix implies a Dirac field $(N^{\prime},N^{c})$, with $N^{\prime}=c_{\alpha}N+s_{\alpha}\nu_{3}$, of mass $m_{N}=\sqrt{M^{2}+m^{2}}$ plus three massless neutrinos. The (massless) neutrino $\nu^{\prime}_{3}=-s_{\alpha}N+c_{\alpha}\nu_{3}$ has now a small component $s_{\alpha}=m/\sqrt{M^{2}+m^{2}}$ along the sterile flavor $N$.
In order to get non-zero masses we need to break $L$. In particular, we can add small Majorana masses ($\mu_{1,2}\ll M$, $\Delta L=2$) for the two heavy fields and suppressed Yukawas ($\tilde{y}_{\nu}\ll y_{\nu}$, $\Delta L=1$) for $N$, implying the matrix
${\cal M}=\begin{pmatrix}&&&0&0\\ &0&&\mu_{3}&0\\ &&&\mu^{\prime}_{3}&m\\ 0&\mu_{3}&\mu^{\prime}_{3}&\mu_{1}&M\\ 0&0&m&M&\mu_{2}\end{pmatrix}.$ (3)
The mass $\mu_{1}$ increases in one unit the rank of ${\cal M}$ and defines an inverse seesaw with $m_{\nu_{3}}\approx\mu_{1}(m/M)^{2}$ [5]. The term $\mu_{2}$ does not give mass to a second standard neutrino, it just breaks the degeneracy between the two (now Majorana) heavy neutrinos. The term $\mu_{3}=\tilde{y}_{\nu}v/\sqrt{2}$ is then necessary to give a mass $m_{\nu_{2}}\approx\mu_{3}^{2}/\mu_{1}$ to $\nu^{\prime}_{2}$, with $\nu^{\prime}_{1}$ staying massless. Alternatively, $\mu_{3}$ could be forbidden by a discrete symmetry and $m_{\nu_{2}}$ obtained by adding a second heavy neutrino pair (i.e., an inverse seesaw for the two massive neutrino families).
[…]
The possibility of two Majorana HNLs at the TeV scale with sizeable heavy-light mixings (e.g., $m=10$ GeV, $\mu_{1,2}\approx 1$ TeV and $M=0$) requires a fine tuning ($m^{2}/\mu_{2}+\mu^{\prime 2}_{3}/\mu_{1}\approx 10^{-8}m^{2}/\mu_{2}$) that is not stable under radiative corrections. If one of the HNLs is significantly heavier than the other one, the cancelation forces its mixing to be smaller, and the model reduces to a single Majorana HNL mixed with a combination of the standard neutrinos. Therefore, although throughout our analysis we will assume a quasi-Dirac HNL $N$, we will also discuss this fine tuned Majorana case.
The heavy-light mixing $s_{\alpha}$ defines then the couplings of $\nu_{\ell}$ and $N$ to the $W^{\pm}$ and $Z$ bosons (we drop the prime to denote mass eigenstates). In particular, $\nu_{\ell}$ will see its gauge couplings reduced by a factor of $c_{\alpha}$ whereas $N$ will now couple to the $W$ with a strength proportional to $V_{\ell N}\approx s_{\alpha}$. The model also implies heavy–light couplings to the $Z$ boson ($\propto s_{\alpha}c_{\alpha}$) and to the Higgs boson ($y_{\nu}=\sqrt{2}s_{\alpha}M/v$).
Collider bounds on $|V_{\ell N}|^{2}$ may be obtained at energies below or above $m_{N}$. At lower energies the HNL is not produced and the bounds are based on observables like the universality of weak interactions, precision observables related to $\mu$ decays (notice that the mixing with $\nu_{\mu}$ changes the definition of $G_{\mu}$), the invisible width of the $Z$ boson or one-loop flavor changing processes like $\mu\to e\gamma$. A global fit of these observables sets limits on $|V_{\ell N}|^{2}$ ranging from $10^{-2}$ to $10^{-4}$, being the most stringent ones for $\ell=\mu$ [6, 7, 8].
Figure 1: Dominant diagrams in $N$ searches at CMS (Dirac or Majorana) and L3.
Here we will discuss higher energy processes with the direct production of the HNL. The current collider limits for $m_{N}>M_{W}$ have been obtained by CMS in the trilepton channel at the LHC [9, 10] and by L3 at LEP [11] (see the dominant diagrams in Fig. 1), in this second case only for $\ell=e$. We will focus on the trilepton channel. We will show that, although the longitudinal momentum of the $\nu$ escapes detection, the mass $m_{N}$ of the HNL can be reconstructed. We will also comment on the possible effects introduced by a HNL in Higgs searches at the LHC. In particular, notice (see Fig. 2) that the dilepton plus missing $p_{\rm T}$ signal from $h\to WW^{*}$ observed at CMS coincides with the one from $h,Z\to\nu N$, although the kinematics in each process is obviously different.
Figure 2: Higgs to $W^{+}W^{-}$ in the SM and possible HNL contributions to dilepton plus $p^{\rm miss}_{\rm T}$.
2 HNL at the LHC
To study a HNL at the LHC we first produce an UFO file [12] with FeynRules [13] that includes the new interactions, then we generate signal and background events with MadGraph5 [14] + Pythia [15], and finally we analyze the results with routines based on FasJet [16], ROOT [17] and HepMC [18].
Let us consider a $90$, $110$ or $130$ GeV HNL with, in all cases, a mixing $|V_{\mu N}|^{2}=10^{-3}$ with the muon flavor. For 138 fb-1 at 13 TeV, our simulation estimates, respectively, 1021, 317 or 153 HNL events giving two opposite sign muons, an electron and $\vec{p}_{\rm T}^{\rm\,miss}$:
$pp\to\mu^{\pm}N\to\mu^{\pm}\mu^{\mp}W^{\pm}\to\mu^{\pm}\mu^{\mp}e^{\pm}\nu\,.$ (4)
In the simulation we have included the events where the final electron comes from the leptonic decay of a tau lepton (i.e., the $\ell^{\prime+}$ in Fig. 1 may be an $e^{+}$ or a $\tau^{+}$ decaying into $\bar{\nu}_{\tau}\,e^{+}\nu_{e}$). We then apply the cuts defined in [10], namely, the three leptons must be isolated and with a $p_{\rm T}>10$ GeV (highest $p_{\rm T}$ above 55 GeV), a pseudorapidity $|\eta|<2.4\,(2.5)$ for the muons (electron), and an invariant mass $m_{\mu\mu}>5$ GeV with $|m_{\mu\mu}-m_{Z}|>15$ GeV and $|m_{\mu\mu e}-m_{Z}|>15$ GeV. We find that 20 events (just 2% of the initial HNL sample) pass the cuts for $m_{N}=90$ GeV, 17 events for $m_{N}=110$ GeV and 14 for $m_{N}=130$ GeV. The low efficiency of the cuts suggests that the CMS analysis has been optimized for larger HNL massesFor $m_{N}=90$ GeV, in particular, we find that a reduction from 10 to 5 GeV in the minimum $p_{T}$ of the least energetic lepton increases the signal by a factor of 1.42 and the background by a factor of 1.12)..
The background, in turn, is dominated by $WZ$ and $ZZ$ production: we obtain that the cuts are satisfied by 285 events, which define around 50% of a total background yield [10] (these 285 events do not include leptons from hadron decays nor photons and jets misidentified as leptons).
The momenta of the two muons and the electron in the HNL events can be measured, whereas the transverse momentum $\vec{p}^{\,\nu}_{\rm T}$ of the neutrino can be obtained from $\vec{p}_{\rm T}^{\rm\,miss}$. In addition, we notice that for $m_{N}>m_{W}$ the $W$ boson decaying into electron plus neutrino is on shell. This can be used to deduce the longitudinal momentum $p_{\rm L}^{\nu}$ of the neutrinoFor $\ell^{\prime}=\tau$ (see Fig. 1) $p_{\rm L}^{\nu}$ refers to the total momentum of the three final neutrinos.: it is solution to the quadratic equation
$(p^{e}_{\rm T})^{2}\,(p_{\rm L}^{\nu})^{2}-p^{e}_{\rm L}\left(m_{W}^{2}+2\,\vec{p}^{\,e}_{\rm T}\cdot\vec{p}^{\,\nu}_{\rm T}\right)p_{\rm L}^{\nu}-\left({m_{W}^{4}\over 4}+m_{W}^{2}\,\vec{p}^{\,e}_{\rm T}\cdot\vec{p}^{\,\nu}_{\rm T}+(\vec{p}^{\,e}_{\rm T}\cdot\vec{p}^{\,\nu}_{\rm T})^{2}-(p^{e})^{2}\,(p^{\nu}_{\rm T})^{2}\right)=0\,.$ (5)
With the complete momenta of the neutrino and of the electron and the muon of opposite charge (i.e., $e^{\pm}\mu^{\mp}$) we can now reconstruct $m_{N}$,
$m_{N}=\sqrt{\left(E^{\mu}+E^{e}+E^{\nu}\right)^{2}-\left(\vec{p}^{\,\mu}+\vec{p}^{\,e}+\vec{p}^{\,\nu}\right)^{2}}\,.$ (6)
Only one of the two solutions for $p_{L}^{\nu}$ in Eq (5) provides the right reconstruction of $m_{N}$, and we find the lower one to be the best choice. In particular, when the two solutions are significantly different ($m_{2}>1.1\,m_{1}$) this is the right choice 96% of the times, whereas for similar solutions the prescription is successful 84% of the times.
Figure 3: Mass reconstruction in $pp\to\mu^{\pm}\mu^{\mp}e^{\pm}\nu$ (left) and $pp\to\mu^{\pm}\mu^{\mp}\mu^{\pm}\nu$ (right) at 13 TeV, 138 fb-1 for $m_{N}=(90,110,130)$ GeV with $|V_{\mu N}|^{2}=10^{-3}$ and diboson background.
In Fig. 3 we plot the distribution of the reconstructed $m_{N}$ for the 20, 17 and 14 signal events that correspond to the three HNL masses together with the 285 background events (these distributions have been obtained with a normalized sample 20 times larger). It is apparent that the peak at $m_{N}=90$ or $110$ GeV would be significant enough to exclude the mixing $|V_{\mu N}|^{2}=10^{-3}$ that we are considering.
We can increase the statistics if we include the possibility that the $W$ boson decays into a muon,
$pp\to\mu^{\pm}N\to\mu^{\pm}\mu^{\mp}W^{\pm}\to\mu^{\pm}\mu^{\mp}\mu^{\pm}\nu\,.$ (7)
In this case, the muon needed to reconstruct $m_{N}$ appears together with a second muon with the same charge. We notice, however, that the production of a HNL (see Fig. 1) is favored by the collision of a valence quark and a see antiquark in the initial protons. This tends to give a larger $|p_{\rm L}|$ to the muon in the primary vertex and, together with the cut $p^{\rm miss}_{\rm T}>55$ GeV, implies that $m_{N}$ must be reconstructed with the muon of lower energy. The Monte Carlo simulation shows that this prescription gives the right choice 78% of the times for $m_{N}=90$ GeV or 87% for $m_{N}=130$ GeV.
Therefore we consider as well this trimuon signal. Again, we find that only 2% of the HNL events pass the cuts for $m_{N}=90$ GeV or 6% at $m_{N}=130$ GeV. The reconstruction of $m_{N}$ is done with the criteria described above but changing the electron by the less energetic muon among the two with same charge. The event distribution for this observable is shown in Fig. 3-right. Combined with the dimuon plus electron events and normalizing the background by a factor of 2§§§As shown in [10], nonprompt leptons and missidentified jets and photons account approximately for $50\%$ of the background., we estimate that at the current luminosity the bounds on $|V_{\mu N}|^{2}$ at CMS and ATLAS could be around $4.0\times 10^{-4}$ for $m_{N}=90$ GeV or $8.4\times 10^{-4}$ for $m_{N}=130$ GeV. These limits are set at 95% confidence level using the $CL_{s}$ prescription [19]. An analogous analysis for $pp\to e^{\pm}N\to e^{\pm}e^{\mp}\ell^{\prime\pm}\nu$ (in Fig. 4) would imply very similar bounds for the mixing with the electron flavor: $|V_{eN}|^{2}<3.8\times 10^{-4}$ or $8.0\times 10^{-4}$ for a HNL mass of $90$ or $130$ GeV, respectively.
Figure 4: Mass reconstruction in $pp\to e^{\pm}e^{\mp}\mu^{\pm}\nu$ (left) and $pp\to e^{\pm}e^{\mp}e^{\pm}\nu$ (right) at 13 TeV, 138 fb-1 for $m_{N}=(90,110,130)$ GeV with $|V_{eN}|^{2}=10^{-3}$ and diboson background.
3 Majoranas and other production channels
Figure 5: Mass reconstruction in $pp\to\mu^{\pm}\mu^{\pm}e^{\mp}\nu$ for a Majorana mass $m_{N}=(90,110,130)$ GeV with $|V_{\mu N}|^{2}=10^{-3}$ at 13 TeV aat 13 TeV, 138 fb-1.
We can also reconstruct $m_{N}$ for a heavy Majorana singlet. This case implies $L$-violating processes with $p_{\rm T}^{\rm miss}$, two muons (or electrons) with the same charge plus an electron (or muon) of opposite charge (see Fig. 1), e.g.,
$pp\to W^{+}\to\mu^{+}N\to\mu^{+}\mu^{+}\,W^{-}\to\mu^{+}\mu^{+}e^{-}\,\bar{\nu}_{e}\,.$ (8)
To reconstruct $m_{N}$ we need to choose among the two same sign muons. As explained before, the muon from $W^{+}\to\mu^{+}N$ tends to be the most energetic one. Here, in addition, the chirality flip in $N$ favors that the $\mu^{+}$ from $N\to\mu^{+}W^{-}$ is emitted backwards [20], which also reduces its energy: 89% of the times the reconstruction with the less energetic muon provides the right value for $m_{N}=110$ GeV. Again, the efficiency of the cuts seems low: 19 out of 511 HNL events pass the cuts for $m_{N}=90$ GeV or 10 out of 83 for $m_{N}=130$ GeV. Fig. 5 shows, however, that the reconstruction of $m_{N}$ works and, together with a total background 50% smaller than in the Dirac case, our results suggest similar limits. For example, at $m_{N}=90$ GeV we estimate
$|V_{eN}|^{2}<4.0\times 10^{-4}\,;\hskip 14.22636pt|V_{\mu N}|^{2}<4.2\times 10^{-4}\,.$ (9)
Finally, we would like to comment on the reconstruction of the HNL mass also in other channels. In particular, Higgs searches $h\to WW^{*}\to\mu^{+}\nu\,e^{-}\bar{\nu}$ (dilepton of different flavor plus missing $p_{\rm T}$) [21, 22, 23, 24, 25] with slightly different cuts could be sensitive to the presence of the HNL: $Z(h)\to\nu N\to\nu\,\mu^{+}W^{-}\to\nu\,\ell^{+}\bar{\nu}\,e^{-}$.
Take $m_{N}=85$ GeV (the argument applies to any masses between $m_{W}$ and around $100$ GeV) and $|V_{\mu N}|^{2}=10^{-3}$. Before the cuts described in [26], we estimate 9,950 HNL events versus 11,320 Higgs events. We notice that the low value of $m_{N}$ strongly favors that the s-channel $Z$ boson in Fig. 2 is near the mass shell, implying that the neutrino from $Z\to\nu N$ will carry little $p_{\rm T}$. We can then assume that most of the missing $p_{\rm T}$ is carried by the second neutrino and, as before, use that the parent $W$ is on shell to find $p^{\nu}_{L}$ and reconstruct $m_{N}$. We find, however, that with the cuts optimized for Higgs searches very few $N$ events are selected: the ones used in [26] keep only 0.7% of the 9,950 events, versus 11% of the 11,320 Higgs events. The basic reason is that the charged lepton from $N\to\mu^{+}W^{-}$ tends to have a $p_{\rm T}$ below the 15 GeV required to the subleading lepton. If we relax this minimum $p_{\rm T}$ to 5 GeV and impose that the electron is more energetic than the muon we would keep 1.1% of the $N$ sample (106 events), while the fraction of $h\to WW^{*}$ events passing the cuts would be reduced to 4.4% (500 Higgs events). These frequencies underline the possible complementarity of Higgs physics and $N$ searches in some mass region.
4 Discussion
In any search for a new particle it is critical to find the optimal kinematical variable that discriminates between signal and background. Here we have discussed how to reconstruct the mass of a HNL in the trilepton plus $p_{\rm T}^{\rm miss}$ channel. Our Monte Carlo simulation shows a clear peak in this variable, that would translate into better bounds on the heavy-light mixing $|V_{\ell N}|^{2}$. In particular, a recent CMS analysis [10] uses a combination of observables optimized with machine learning techniques that distributes both signal and background in 20 bins. Our results clearly indicate that $m_{N}$ would be a better choice in this case.
In another recent search for a light ($m_{N}<m_{W}$) HNL [27], ATLAS has reconstructed $m_{N}$ using that the $s$-channel $W$ in $pp\to W^{+}\to\ell^{+}N$ (see Fig. 1) is near the mass shell. Since the $W$ in $N\to\ell^{-}W^{+}$ is exactly on shell, our analysis provides a more accurate reconstruction in the complementary regime with $m_{N}>m_{W}$. We have also shown that a similar reconstruction could work in the dilepton plus $p_{\rm T}^{\rm miss}$ channel by changing the cuts currently being used in Higgs to $WW^{*}$ searches.
In summary, although it is a possibility not explored at the LHC yet, we show that the search for peaks associated to a HNL in $pp$ collisions could probe mixings $|V_{\ell N}|^{2}$ that are not excluded by PMNS unitarity bounds.
Acknowledgments
We would like to thank Mikael Chala and José Santiago for discussions. This work has been supported by the Spanish Ministry of Science, Innovation and Universities MICIU/AEI/ 10.13039/501100011033/ (grants PID2022-14044NB-C21 and PID2022-139466NB-C22) and by Junta de Andalucía (FQM 101).