### v4

# Reconstructing a Heavy Neutral Lepton at the LHC

Abstract
Heavy lepton singlets $N$ slightly mixed with a standard neutrino $\nu_{\ell}$ are usually searched for at the LHC in the trilepton plus $p_{\rm T}^{\rm miss}$ channel: $pp\to W^{+}\to\ell^{+}N$ with $N\to\ell^{-}W^{+}\to\ell^{-}\ell^{\prime+}\nu$. We show that, although the longitudinal momentum of the final $\nu$ escapes detection, the mass of the heavy lepton can be reconstructed. While this possibility has not been considered in recent LHC searches, we find that the search for a mass peak could systematically improve the current collider bounds on the mixing $|V_{\ell N}|^{2}$ for any mass $m_{N}\geq M_{W}$.
Reconstructing a Heavy Neutral Lepton at the LHC
Pablo de la Torre, Manuel Masip, Fuensanta Vilches
Departamento de Física Teórica y del Cosmos
Universidad de Granada, E-18071 Granada, Spain
pdelatorre@ugr.es      masip@ugr.es      fuenvilches@ugr.es
1 Introduction
Heavy neutral leptons (HNL) provide the simplest UV completion of the Standard Model (SM) that is able to generate the dim-5 Weinberg operator [1], which arguably makes them the most motivated scenario for BSM physics [2, 3, 4, 5, 6, 7, 8]. Let us briefly discuss a minimal setup (see also the models in [9, 10]) using 2-component spinors.
To explain neutrino masses we need at least two bi-spinors $(N,N^{c})$ of opposite lepton number $L=\pm 1$ combined into a Dirac field of mass $M$. Assuming no extra Higgses, lepton-number conservation and only dim $\leq 4$ operators we have
$-{\cal L}\supset M\,NN^{c}+y_{\nu}\,HLN^{c}+{\rm h.c.}\,,$ (1)
where $H=(h^{+}\;h^{0})$, $L=(\nu\;\ell)$ and we have redefined the 3 lepton doublets to obtain the flavor combination in the Yukawa. After electroweak (EW) symmetry breaking the neutrino mass matrix reads
${\cal M}=\begin{pmatrix}&&&\cdot&0\\ &0&&\cdot&0\\ &&&\cdot&m\\ \cdot&\cdot&\cdot&\cdot&M\\ 0&0&m&M&\cdot\end{pmatrix},$ (2)
where $m=y_{\nu}v/\sqrt{2}$ and the dots indicate terms forbidden by $L$. This rank-2 matrix implies a Dirac field $(N^{\prime},N^{c})$, with $N^{\prime}=c_{\alpha}N+s_{\alpha}\nu_{3}$, of mass $m_{N}=\sqrt{M^{2}+m^{2}}$ plus three massless neutrinos. The neutrino $\nu^{\prime}_{3}=-s_{\alpha}N+c_{\alpha}\nu_{3}$ is massless but has now a small component $s_{\alpha}=m/\sqrt{M^{2}+m^{2}}$ along the sterile flavor $N$.
In order to get non-zero masses we need to break $L$: we can add small Majorana masses ($\mu_{1,2}\ll M$, $\Delta L=2$) for the two heavy modes and suppressed Yukawas ($\tilde{y}_{\nu}\ll y_{\nu}$, $\Delta L=1$) for $N$, implying
${\cal M}=\begin{pmatrix}&&&0&0\\ &0&&\mu_{3}&0\\ &&&\mu^{\prime}_{3}&m\\ 0&\mu_{3}&\mu^{\prime}_{3}&\mu_{1}&M\\ 0&0&m&M&\mu_{2}\end{pmatrix}.$ (3)
The mass $\mu_{1}$ increases in one unit the rank of ${\cal M}$ and defines an inverse seesaw with $m_{\nu_{3}}\approx\mu_{1}(m/M)^{2}$ [11]. The term $\mu_{2}$ does not give mass to a second standard neutrino, it just breaks the degeneracy between the two (now Majorana) heavy neutrinos. The term $\mu_{3}=\tilde{y}_{\nu}v/\sqrt{2}$ is then necessary to give a mass $m_{\nu_{2}}\approx\mu_{3}^{2}/\mu_{1}$ to $\nu^{\prime}_{2}$, with $\nu^{\prime}_{1}$ staying massless. Alternatively, $\mu_{3}$ could be forbidden by a discrete symmetry and $m_{\nu_{2}}$ obtained by adding a second heavy neutrino pair (i.e., an inverse seesaw for the two massive neutrino families, with the third neutrino massless).
The mass parameters $M$ and $\mu_{1,2}$ above are not EW, and the matrix in Eq. (3) may accommodate any values between $10^{-3}$ and $10^{15}$ GeV: any HNL mass seems equally natural, and only the data may have a say about nature’s choice. Masses larger than the TeV, however, imply a very small mixing $s_{\alpha}$ and thus a decoupled HNL. In particular, the usual seesaw mechanism is obtained if the three mass parameters and the Yukawa couplings are all unsuppressed, e.g., $M,\mu_{1,2}\approx 10^{10}$ GeV and $m,\mu_{3}\approx 1$ GeV.
The possibility of two Majorana HNLs at the TeV scale with sizeable heavy-light mixings (e.g., $m=10$ GeV, $\mu_{1,2}\approx 1$ TeV and $M=0$) requires a fine tuning ($m^{2}/\mu_{2}+\mu^{\prime 2}_{3}/\mu_{1}\approx 10^{-8}m^{2}/\mu_{2}$) that is not stable under radiative corrections. If one of the HNLs is significantly heavier than the other one, the cancelation forces its mixing to be smaller and the model reduces to a single Majorana HNL mixed with a combination of the standard neutrinos. Therefore, although throughout our analysis we will assume a quasi-Dirac HNL $N$, we will also comment on this fine tuned Majorana case.
The heavy-light mixing $s_{\alpha}$ defines then the couplings of $\nu_{\ell}$ and $N$ to the $W^{\pm}$ and $Z$ bosons (we drop the prime to denote mass eigenstates). In particular, $\nu_{\ell}$ will see its gauge couplings reduced by a factor of $c_{\alpha}$ whereas $N$ will now couple to the $W$ with a strength proportional to $V_{\ell N}\approx s_{\alpha}$. The model also implies heavy–light couplings both to the $Z$ boson ($\propto s_{\alpha}c_{\alpha}$) and to the Higgs boson ($y_{\nu}=\sqrt{2}s_{\alpha}M/v$). We will consider the case where $N$ couples to a single generation of SM neutrinos [12, 8].
Collider bounds on $|V_{\ell N}|^{2}$ may be obtained at energies below or above $m_{N}$. At lower energies the HNL is not produced and the bounds are based on observables like the universality of weak interactions, precision observables related to $\mu$ decays (the mixing with $\nu_{\mu}$ changes the definition of $G_{F}$), the invisible width of the $Z$ boson or one-loop flavor changing processes like $\mu\to e\gamma$. A global fit of these observables sets PMNS unitarity limits on $|V_{\ell N}|^{2}$ ranging from $10^{-2}$ to $10^{-4}$, being the most stringent ones for $\ell=\mu$ [13, 14, 15, 16].
Figure 1: Dominant diagrams in $N$ searches at CMS (Dirac or Majorana) and L3.
Here we will discuss higher energy processes with the direct production of the HNL. The signatures of such models at colliders have been studied extensively [17, 18, 19, 20, 21, 22, 23, 24, 25]. The current bounds on $|V_{\ell N}|^{2}$ for $m_{N}>M_{W}$ have been obtained by CMS in the trilepton channel at the LHC [26, 27] and by L3 at LEP [28], in this second case only for $\ell=e$. We will focus on the trilepton channel, a process dominated by the charged-current Drell–Yan $q\bar{q}^{\prime}\to W^{\pm}\to Nl^{\pm}$ [29, 30] (see Fig. 1). In our analysis we will also include vector boson fusion $q\gamma\to Nl^{\pm}q^{\prime}$ [31, 32, 33], a process giving a significant contributions for masses $m_{N}\geq 500$ GeV. We will show that, although the longitudinal momentum of the $\nu$ escapes detection, the mass $m_{N}$ of the HNL can be reconstructed.
We will comment as well on the possible effects introduced by an HNL in Higgs searches at the LHC. In particular, notice (see Fig. 2) that the dilepton plus missing $p_{\rm T}$ signal from $h\to WW^{*}$ observed at CMS coincides with the one from $h,Z\to\nu N$, although the kinematics in each process is obviously different.
Figure 2: Higgs to $W^{+}W^{-}$ in the SM and possible HNL contributions to dilepton plus $p^{\rm miss}_{\rm T}$.
2 HNL at the LHC
To estimate the possible signal of an HNL at the LHC, we first produce a next-to-leading-order (NLO) UFO file [34] with the new vertices using the Mathematica packages FeynRules [35], NLOCT [36] and FeynArts [37]. Then, we generate signal and background events at NLO using MadGraph5_aMC@NLO v2.9.7 [38] interfaced with Pythia v8.310 [39], MadSpin [40] and Delphes v3.2.0 [41]. In Delphes we have used the CMS simulation card, including a photon conversion module based on the characteristics of the detector [42, 43, 44, 45]. For the parton distribution functions (PDFs), we use NNPDF3.1 [46] in all the cases except for vector boson fusion signal process, where we use NNPDF31.luxQED [47, 48, 49] (both as implemented in lhapdf [50]). For the merging between jets from matrix elements and from parton showers, we have used the FxFx matching scheme [51]. Finally, we have analyzed the results with custom routines based on FasJet [52], ROOT [53] and HepMC [54].
After we generate the events, we apply the cuts defined in [27] for the high-mass regime, with at least one opposite-sign same-flavour (OSSF) lepton pair and
  1. 1.
    The three leptons must be isolated as defined in [55], with $p_{\rm T}(\ell_{1})>55$ GeV, $p_{\rm T}(\ell_{2})>15$ GeV, $p_{\rm T}(\ell_{3})>10$ GeV.
  2. 2.
    No $b$ jets.
  3. 3.
    Any OSSF pair must satisfy $m(\ell^{+}\ell^{-})>5$ GeV and $\left|m(\ell^{+}\ell^{-})-m_{Z}\right|>15$ GeV.
  4. 4.
    The trilepton invariant mass is $\left|m(3\ell)-m_{Z}\right|>15$ GeV.
To illustrate our results, let us consider three benchmarks masses $m_{N}$ in the 90909090170170170170 GeV range and two more in the higher mass region 400400400400600600600600 GeV. In all the cases we assume a mixing $|V_{\mu N}|^{2}=10^{-2}$ with the muon flavor (the cross sections are proportional to $|V_{\mu N}|^{2}$) and we take a luminosity of 138 fb-1 at 13 TeV. We first focus on the channel
$pp\to\mu^{\pm}N\to\mu^{\pm}\mu^{\mp}W^{\pm}\to\mu^{\pm}\mu^{\mp}e^{\pm}\nu\,.$ (4)
[…]
The background, in turn, is made up of $WZ$ (37%), $ZZ$ (1%), $Z\gamma$ (38%), nonprompt ($Z$+jets, $t\bar{t}$+jets, 20%) and other ($WWW$, $WWZ$, $ZZ$, $t\bar{t}W$, $t\bar{t}Z$, $t\bar{t}\gamma$, 4%). We estimate 54840 events that are reduced to 1259 background events after the cuts. These frequencies are consistent (within a 20%) with the ones used in [27].
$m_{N}$ [GeV] 90 130 170 400 600
$\mu^{\pm}\mu^{\mp}e^{\pm}$ events 858 297 101 4.01 1.78
events after cuts 148 119 72.5 3.87 1.73
(17.2%) (40.0%) (71.8%) (96.5%) (97.2%)
Table 1: Signal events in the channel $\mu^{\pm}\mu^{\mp}e^{\pm}\nu$ before and after cuts (138 fb-1 at 13 TeV) for several HNL masses and $|V_{\mu N}|^{2}=10^{-2}$.
Now, the detector provides the momenta of the two muons and of the electron in all the signal events, whereas the transverse momentum $\vec{p}^{\,\nu}_{\rm T}$ of the neutrino can be obtained from $\vec{p}_{\rm T}^{\rm\,miss}$. In addition, we notice that for $m_{N}>m_{W}$ the $W$ boson decaying into electron plus neutrino is on shell. This can be used to deduce the longitudinal momentum $p_{\rm L}^{\nu}$ of the neutrino (see [56] for an analogous calculation in a different context): it is solution to the quadratic equation
$(p^{e}_{\rm T})^{2}\,(p_{\rm L}^{\nu})^{2}-p^{e}_{\rm L}\left(m_{W}^{2}+2\,\vec{p}^{\,e}_{\rm T}\cdot\vec{p}^{\,\nu}_{\rm T}\right)p_{\rm L}^{\nu}-\left({m_{W}^{4}\over 4}+m_{W}^{2}\,\vec{p}^{\,e}_{\rm T}\cdot\vec{p}^{\,\nu}_{\rm T}+(\vec{p}^{\,e}_{\rm T}\cdot\vec{p}^{\,\nu}_{\rm T})^{2}-(p^{e})^{2}\,(p^{\nu}_{\rm T})^{2}\right)=0\,.$ (5)
With the complete momenta of the neutrino and of the electron and the muon of opposite charge (i.e., $e^{\pm}\mu^{\mp}$) we can now reconstruct $m_{N}$,
$m_{N}=\sqrt{\left(E^{\mu}+E^{e}+E^{\nu}\right)^{2}-\left(\vec{p}^{\,\mu}+\vec{p}^{\,e}+\vec{p}^{\,\nu}\right)^{2}}\,.$ (6)
We find, however, that after the detector simulation 8% of the signal events for $m_{N}=90$ GeV or 13% for 600 GeV have no real solution to Eq (5). This introduces an effective extra cut in the signal that, interestingly, is stronger on the background: 25% of the background events are cut by the requirement of having a solution to Eq (5). For the rest of events the equation implies a two-fold degeneracy, and we find the lower solution to be the best choice 87% of the times.
In Fig. 3 we plot the reconstruction of different values of $m_{N}$ in the low and high mass regions together with the background.
Figure 3: Reconstruction of $m_{N}$ from $pp\to\mu^{\pm}\mu^{\mp}e^{\pm}\nu$ in the low (left) and high (right) mass regions for $|V_{\mu N}|^{2}=0.01$ and $0.1$, respectively, together with the background (from top to bottom, the contributions from $Z\gamma$, $WZ$, $ZZ$ and nonprompt+other separated by dashes).
We can increase the statistics if we include the possibility that the $W$ boson decays into a muon,
$pp\to\mu^{\pm}N\to\mu^{\pm}\mu^{\mp}W^{\pm}\to\mu^{\pm}\mu^{\mp}\mu^{\pm}\nu\,.$ (7)
We plot the number of events before and after cuts in Table 2. This trimuon channel has an estimated background of 31321 events (60% from $WZ$, 36% from $ZZ$, 3% nonprompt and below1% from other) that is reduced to 642 events after cuts.
In this case, the muon needed to reconstruct $m_{N}$ appears together with a second muon with the same charge. We notice, however, that the production of an HNL (see Fig. 1) is favored by the collision of a valence quark and a sea antiquark in the initial protons. This tends to give a larger $p_{\rm T}$ to the muon in the primary vertex and implies that $m_{N}$ must be reconstructed with the muon of lower $p_{\rm T}$. We find that this prescription works 92% of the times for $m_{N}=130$ GeV or 77% for $m_{N}=400$ GeV.
Therefore we consider as well this trimuon channel. The reconstruction is done with the same criteria as before but changing the electron by the muon with less $p_{\rm T}$ among the two with same charge. The distribution for the reconstructed $m_{N}$ is shown in Fig. 3 in the two mass regimes. Combining the $\mu\mu e$ and the $\mu\mu\mu$ channels, we obtain an estimate of the expected LHC limits at 95% confidence level (we use the $CL_{s}$ prescription [57]). We find bounds (see Table 3) that go from $|V_{\mu N}|^{2}<7.8\times 10^{-4}$ for $m_{N}=90$ GeV to $|V_{\mu N}|^{2}<6.2\times 10^{-2}$ for $m_{N}=600$ GeV.
$m_{N}$ [GeV] 90 130 170 400 600
$\mu^{\pm}\mu^{\mp}\mu^{\pm}$ events 981 336 112 4.39 1.90
events after cuts 164 92.3 53.7 4.13 1.85
(16.7%) (27.4%) (47.9%) (94.1%) (97.3%)
Table 2: Signal events in the channel $\mu^{\pm}\mu^{\mp}\mu^{\pm}\nu$ before and after cuts (138 fb-1 at 13 TeV) for several HNL masses and $|V_{\mu N}|^{2}=10^{-2}$.
We can estimate the possible relevance of this observable in current collider searches. In a recent analysis [27], CMS set bounds using boosted decision trees (BDT) trained with several kinematic observables (mainly $\Delta\text{R}\left[\text{min }m(l^{+}l^{-})\right]$, $m_{T}$, $p_{T}(l_{3})$ and $p_{T}^{\text{miss}}$). Our results indicate that this optimized combination of observables provides a less efficient signal-background discrimination than $m_{N}$, especially in the low mass region: $|V_{\mu N}|^{2}<2.2\times 10^{-3}$ (versus $7.8\times 10^{-4}$) for $m_{N}=90$ GeV or $|V_{\mu N}|^{2}<5.6\times 10^{-3}$ (versus $2.6\times 10^{-4}$) for $m_{N}=130$ GeV. Therefore, significantly better bounds could be expected if $m_{N}$ were added to the pool of observables training the BDT.
Figure 4: Reconstruction of $m_{N}$ from $pp\to\mu^{\pm}\mu^{\mp}\mu^{\pm}\nu$ in the low (left) and high (right) mass regimes for $|V_{\mu N}|^{2}=0.01$ and $0.1$, respectively, together with the background (from top to bottom, the contributions from $WZ$, $ZZ$ and nonprompt+other separated by dashes).
$m_{N}$ [GeV] 90 130 170 400 600
$|V_{\mu N}|^{2}$ limit $7.8\times 10^{-4}$ $2.6\times 10^{-3}$ $4.9\times 10^{-3}$ $3.8\times 10^{-2}$ $6.2\times 10^{-2}$
Table 3: Expected bounds on $|V_{\mu N}|^{2}$ by combining the channels $\mu^{\pm}\mu^{\mp}e^{\pm}\nu$ and $\mu^{\pm}\mu^{\mp}\mu^{\pm}\nu$ for 138 fb-1 at 13 TeV.
An analogous analysis can be done for the mixing $|V_{eN}|$,
$pp\to W^{\pm}\to e^{\pm}N\to e^{\pm}e^{\mp}\,W^{\pm}\to e^{\pm}e^{\pm}\ell^{\mp}\,\nu\,.$ (8)
Our results suggests similar bounds in this case. For example, we obtain a 20% smaller signal after cuts at $m_{N}=130$ GeV within a 10% smaller background, implying an expected bound around $|V_{eN}|^{2}<2.9\times 10^{-3}$ for this HNL mass.
3 Majoranas and other production channels
We can also reconstruct $m_{N}$ for a heavy Majorana singlet. We focus in the $L$-violating processes with $p_{\rm T}^{\rm miss}$, two muons (or electrons) with the same charge plus an electron (or muon) of opposite charge (see Fig. 1), e.g.,
$pp\to W^{+}\to\mu^{+}N\to\mu^{+}\mu^{+}\,W^{-}\to\mu^{+}\mu^{+}e^{-}\,\bar{\nu}_{e}\,.$ (9)
To reconstruct $m_{N}$ we need to choose among the two same sign muons. As explained before, the muon from $W^{+}\to\mu^{+}N$ tends to be the one with larger $p_{\rm T}$ Here, in addition, the chirality flip in $N$ favors that the $\mu^{+}$ from $N\to\mu^{+}W^{-}$ is emitted backwards [58], which also reduces its energy: for $m_{N}=130$ GeV, 89% of the times the reconstruction with the muon that has lower $p_{\rm T}$ provides the right value.
Although we find that the reconstruction of $m_{N}$ works, the number of signal events in the channel $\mu^{+}\mu^{+}e^{-}$ is half the one in $\mu^{+}\mu^{-}e^{+}$ for the Dirac case, plus here the trimuon channel does not provide an efficient reconstruction of $m_{N}$. This, together with a background that is now only composed by $t\bar{t}$+jets, $WWW$ and $t\bar{t}W$, imply similar limits: $|V_{\mu N}|^{2}<2.1\times 10^{-3}$ and $|V_{eN}|^{2}<2.5\times 10^{-3}$ for $m_{N}=130$ GeV.
Finally, we would like to comment on the reconstruction of the HNL mass also in other channels. In particular, we would like to argue that Higgs searches $h\to WW^{*}\to\mu^{+}\nu\,e^{-}\bar{\nu}$ (dilepton of different flavor plus missing $p_{\rm T}$) [59, 60, 61, 62, 63] with slightly different cuts could be sensitive to the presence of the HNL:
$pp\to Z(h)\to\nu N\to\nu\,\mu^{+}W^{-}\to\nu\,\ell^{+}\bar{\nu}\,e^{-}.$ (10)
Take $m_{N}=85$ GeV (the argument applies to any masses between $m_{W}$ and around $100$ GeV) and $|V_{\mu N}|^{2}=10^{-3}$. Before the cuts described in [64], we estimate 9950 HNL events versus 11320 Higgs events. We notice that the low value of $m_{N}$ strongly favors that the s-channel $Z$ boson in Fig. 2 is near the mass shell, implying that the neutrino from $Z\to\nu N$ will carry little $p_{\rm T}$. We can then assume that most of the missing $p_{\rm T}$ is carried by the second neutrino and, as before, use that the parent $W$ is on shell to find $p^{\nu}_{L}$ and reconstruct $m_{N}$. We find, however, that with the cuts optimized for Higgs searches very few $N$ events are selected: the cuts used in [64] keep only 0.7% of the 9950 events, versus 11% of the 11320 Higgs events. The basic reason is that the charged lepton from $N\to\mu^{+}W^{-}$ tends to have a $p_{\rm T}$ below the 15 GeV required to the subleading lepton. If we relax this minimum $p_{\rm T}$ to 5 GeV and impose that the electron is more energetic than the muon we would keep 1.1% of the initial $N$ sample (106 events), while the fraction of $h\to WW^{*}$ events passing the cuts would be reduced to 4.4% (500 Higgs events). These estimates suggest a possible complementarity of Higgs physics and $N$ searches in this mass region.
4 Discussion
[…]
Here we have discussed the reconstruction of $m_{N}$ in the trilepton plus neutrino channel and how to solve the two-fold degeneracy in the quadratic equation that provides the longitudinal momentum of the neutrino. We find that this equation has no solution in a small fraction (around 10%) of signal events. This is mainly due to the smearing effect introduced by the detector, and also to the different kinematics in signal events where the third lepton comes from a tau lepton. We find interesting that the background includes a much larger fraction (around 25%) of events with no solution to the equation, since this could be used as an extra cut also when studying other observables.
In another recent search for a light ($m_{N}<m_{W}$) HNL [65] (see also [66]), ATLAS has reconstructed $m_{N}$ using that the $s$-channel $W$ in $pp\to W^{+}\to\ell^{+}N$ (see Fig. 1) is near the mass shell. Since the $W$ in $N\to\ell^{-}W^{+}$ is exactly on shell, our analysis provides a more accurate reconstruction in the complementary regime with $m_{N}>m_{W}$. We have also shown that a similar reconstruction may work in the dilepton plus $p_{\rm T}^{\rm miss}$ channel by changing the cuts currently being used in Higgs to $WW^{*}$ searches.
In summary, the mass peak discussed here could complement and simplify the search for an HNL in $pp$ collisions and make the details of the calculation less determinant: even if it is not as clean as the ones in Higgs decays, the peak is largely insensitive to normalization uncertainties related to luminosity, overall cross sections, or selection efficiencies. Although it is a possibility not explored at the LHC yet, our results suggest that its search could help to set collider limits on $|V_{\ell N}|^{2}$ competitive with PMNS unitarity bounds.
Acknowledgments
[…]